OpenStudy (anonymous):

how do I find the x-intercepts for y=x^2+8x+15

OpenStudy (agent0smith):

Set y = 0 and factor/solve for x.

OpenStudy (jdoe0001):

\(\large y=x^2+8x+15\implies 0=x^2+8x+15 \quad \begin{array}{cccllll} 0=x^2&+8x&+15\\ &\uparrow &\uparrow \\ &5+3&5\cdot 3 \end{array}\)